OpenStudy (anonymous):

If : \(\color{blue}{\text{2 + 6 + 10 + 14 + 18 + 22 + 26 + 30 + 34 + 38 = 200}}\), Then, \(\color{red}{\text{Add}}\) any \(\color{cyan}{\text{Five}}\) numbers and get \(\color{green}{100}\)..

OpenStudy (anonymous):

5 * 20 = 100 you need to subtract 100, yes? If 20 was negative... that would be -100 So, 200 + (-100) = 100, -100 / 5 = -20, there's five numbers